A 150g bottle has a mass of 300g when filled with water and 310g when filled with an unknown liquid. What is the density of the unknown liquid?
1
Expert's answer
2020-03-16T13:12:30-0400

Solution.

"m(H2O) = 300 \ g - 150 \ g = 150 \ g"

"V(H2O) = \frac{m(H2O)}{\rho(H2O)} = 150 \ cm^3"

"m(X) = 310 \ g - 150 \ g = 160 \ g"

V(H2O) = V(X)

"\rho(X) = \frac{m(X)}{V(X} = 1.07 \ \frac{g}{cm^3}"

Answer:

"\rho(X)= 1.07 \ \frac{g}{cm^3}"